# Simulating stochastic population dynamics: The Linear Noise Approximation can capture non-linear phenomena

### v3

@@ 2.2 System Size Expansion
The time-evolution of the stochastic process $\{{\bm{Y}}(t):t\geq 0\}$ can be described using the random time change representation (RTC) [1]
where, for $j\in\{1,\dots,r\}$, $N_{j}$ are independent, unit Poisson processes corresponding to reaction $R_{j}$. This implies that for a given trajectory ${\bm{Y}}(s)$, $s\in[0,t)$, the random variables $N_{j}\left(\int_{0}^{t}\pi_{j}({\bm{Y}}(s))ds\right)$ are independent and have Poisson distribution with mean $\int_{0}^{t}\pi_{j}({\bm{Y}}(s))ds$.
In pursuit of seeking methods to approximate the process $\{{\bm{Y}}(t):t\geq 0\}$, it is common to introduce a system size parameter $\Omega$. Through setting ${\bm{X}}(t):={\bm{Y}}(t)/\Omega$, one can study the dependence of stochastic fluctuations upon system size. It is sufficient to assume that the rates $\pi_{j}({\bm{Y}}(t))$ depend upon $\Omega$ as $\pi_{j}({\bm{Y}}(t)):=\Omega\rho_{j}({\bm{Y}}(t)/\Omega)$ where $\rho_{j}$ are called macroscopic rates of the system. This is a fairly-relaxed condition, that can be further relaxed to more general relations between the two functions [58, 33].
@@
${\bm{X}}(t+dt)-{\bm{X}}(t)=\sum_{j=1}^{r}{\bm{\nu}}_{j}N_{j}\left(\Omega\rho_{j}({\bm{X}}(t))dt\right)/\Omega,$ (5)
[…]
@@
Letting $dt\to 0$ we derive the macroscopic reaction rate equation (RRE), corresponding to the classical mass-action kinetics. Typical in applied biochemical settings is the need to study the qualitative changes in the dynamics of $\{{\bm{X}}(t):t\geq 0\}$ and $\{{\bm{x}}(t):t\geq 0\}$ when one (or more) of the system parameters, say $\alpha$, is free to vary. To emphasize this dependence on parameter $\alpha\in\mathbb{R}$ we write the RRE as
$\dot{{\bm{x}}}=\frac{d{\bm{x}}}{dt}={\bm{F}}({\bm{x}},\alpha),\quad{\bm{F}}({\bm{x}},\alpha)=\sum_{j=1}^{r}{\bm{\nu}}_{j}\rho_{j}({\bm{x}}(t),\alpha),\qquad{\bm{x}}(0)={\bm{x}}_{0}.$ (6)
and the corresponding Jacobian matrix
@@
Example (SIR model, continued):
@@ 2.3 The Linear Noise Approximation (LNA)
As we show in SI (Section S2) by applying a derivation similar to [1] we get that for sufficiently large values of $\Omega$, the time-evolution of $\{\bm{\xi}(t),t\geq 0\}$ can be described by the linear Stochastic Differential equation (in the Itô sense)
@@
where $\bm{J}_{t}=\bm{J}({\bm{x}}(t))$ is the Jacobian matrix in (7) evaluated at ${\bm{x}}(t)$, ${\bm{B}}_{t}$ an $n$-dimensional Wiener process [64], and
@@
the product of the stoichiometry matrix ${\bm{S}}=[{\bm{\nu}}_{1}\cdots{\bm{\nu}}_{r}]$, and the square root of the diagonal matrix, with main diagonal entries the reaction rates $(\rho_{1}({\bm{x}}(t)),\dots,\rho_{r}({\bm{x}}(t)))$.
This linear SDE in (9) has a solution that, for $0\leq s\leq t<\infty$, can be written as
$\bm{\xi}(t)={\bm{C}}(s,t)\bm{\xi}(s)+{\bm{\eta}}(s,t),\quad{\bm{\eta}}(s,t)\sim\text{MVN}(0,{\bm{D}}(s,t)).$ (11)
@@
[…]
@@
[…]
@@ 2.3.1 Accuracy of the LNA
The LNA accurately approximates the master equation in specific contexts. As discussed in SI (Section S2), the LNA is derived using the Central Limit Theorem, as $\Omega$ tends to $\infty$, as an approximation of the probability distribution function satisfying the master equation. As a result, we expect the LNA to be accurate when $\Omega$ is sufficiently large. The value of $\Omega$ needed to ensure accuracy depends on the characteristics of the system’s dynamics and the length of time internal where the approximation is sought.
For any given finite $\Omega$, if the LNA accurately represents the distribution of the state of a given system at some time-point, it is expected to maintain this accuracy for a subsequent short time interval. The duration for which the LNA remains accurate is influenced by the dynamics of the system and the level of stochasticity present. For instance, it was observed in [41] that for periodic systems with a moderate size $\Omega$, the time interval over which the LNA retains its accuracy is approximately equal to the period of the system.
@@
It is important to note that, for moderate system sizes, the LNA tends to fail in approximating the long-term dynamics of multi-stable and oscillatory systems (see [50, 5, 27, 56, 41] and Figure 2(C-D)).
Figure 2: LNA method summary. A. LNA is a stochastic approximation of the stochastic population dynamics that can be exactly simulated using the SSA. LNA uses the numerical solutions of the deterministic RRE and a SDE that can be solved exactly. The SIR population dynamics model is used here as an example. B. Comparison between SSA and LNA on the SIR population dynamics in terms of simulated $99\%$ Confidence Intervals (CI) shows astonishing agreement. C. In contrast, same comparisons between the LNA and SSA in one of the three species of the oscillatory system in [63] show significant discrepancies. D. Significant discrepancies are also observed for the same comparisons on the bistable Genetic Toggle-Switch system in [12].
2.3.2 Phase
Important in diagnosing this issue, and indeed the rest of this paper, is the idea of phase. For a given reaction network, we can qualitatively examine the dynamics of its paths and identify different stages of their evolution. For example, a phase could correspond to a specific species’ peak or trough or intermediate point in the evolution of an oscillatory system or a point leading to one of the equilibria of a multi-stable system. For a stochastic path $\{{\bm{X}}(t):t\geq 0\}$ and its deterministic path $\{{\bm{x}}(t):t\geq 0\}$ we are able to isolate the points ${\bm{X}}(s)$ and ${\bm{x}}(s)$ at any time $s\geq 0$ on their path and qualitatively say what stages ${\bm{X}}(s)$ and ${\bm{x}}(s)$ are in with respect to the evolution of the entire path. We call these stages the phase of ${\bm{X}}(s)$ and ${\bm{x}}(s)$ respectively. Note that due to stochasticity, the phase of ${\bm{X}}(s)$ and the phase of ${\bm{x}}(s)$ can be different, particularly for large $s$.
We also note that the consideration of phase is only useful for non-linear systems, since for linear systems, all trajectories lead to the same nearby equilibrium as $t\to\infty$ (or $t\to$-$\infty$). For networks with non-linear dynamics (e.g. oscillatory or multi-stable), once the LNA becomes inaccurate at some time-point $\tau$, the phase of ${\bm{X}}(s)$, for $s\geq\tau$, becomes out of sync with the phase of ${\bm{x}}(s)$ causing $\bm{\xi}(s)$ to be too large for the LNA to make accurate predictions, even for short time transitions. That is, for times $s\geq\tau$, there will be no agreement between the LNA and SSA paths, at least for the next time interval. We call this phenomenon a phase drift.
@@ 2.3.2 Phase
Therefore, initialising the LNA to the correct phase is a necessary condition for the LNA to be accurate. This leads us to speculate, if one can continually initialise the LNA with the correct phase so that the phase of the stochastic and deterministic trajectory are in agreement, and $\bm{\xi}(s)$ remains sufficiently small for all $s\geq 0$, then the LNA will remain accurate at all times and produce paths agreeing to those from the SSA. We argue that, for some classes of dynamical systems, this is indeed the case. For instance, it is shown in [41] that this is true for systems with RRE that has an attractive limit cycle solution and a range of moderate system sizes. We adopt the term used in [41] to name this modelling technique, called Phase-Corrected Linear Noise Approximation (pcLNA).
2.4 The phase corrected Linear Noise Approximation (pcLNA)
@@
Here ${G}:\mathbb{R}^{n}\to\gamma$, $G({\bm{X}}(t))={\bm{x}}(s)$, $s\geq 0$, is the phase correction function that maps any stochastic state ${\bm{X}}(t)\in\mathbb{R}^{n}$ to a point, say ${\bm{x}}(s)\in\gamma$, lying on a solution $\gamma=\{{\bm{x}}(t):t\geq 0\}$ of the RRE in (6) that has the correct phase $s\geq 0$ for the current state ${\bm{X}}(t)$. The term $\bm{\kappa}(t)$ denotes the deviation of ${\bm{X}}(t)$ away from ${\bm{x}}(s)={G}({\bm{X}}(t))$. We use this ansatz to simulate stochastic trajectories as described in Algorithm 1.
@@
a. Initial conditions: $s_{0}=0=t_{0}$, ${\bm{x}}(0)$, ${\bm{X}}(0)$, Parameters: $\Delta t$, $\Omega$.
b. A solution $\gamma=\{{\bm{x}}(t):t\geq 0\}$ of a RRE as in (6) and the corresponding solutions $\bm{C}$} and $\bm{D}$} of (12), (13), respectively.
c. The phase correction map $G:\mathbb{R}^{n}\to\gamma$
Steps: Take $\bm{\kappa}(s_{0})=\Omega^{1/2}({\bm{X}}(0)-{\bm{x}}(0))$ and for $i=1,2,\dots$,
@@
  2. 2.
  3. 3.
    set $s_{i}$ such that ${G}({\bm{X}}(t_{i}))={\bm{x}}(s_{i})$
  4. 4.
    set $\bm{\kappa}(s_{i})=\Omega^{1/2}({\bm{X}}(t_{i})-{\bm{x}}(s_{i}))$.
Outputs: $\{{\bm{X}}(t_{i}):i=0,1,2,\dots\}$.
[…]
Figure 3: General pcLNA algorithm steps. pcLNA proceeds with standard LNA steps interrupted by phase corrections, ${\bm{x}}(t_{i})\to G({\bm{X}}(t_{i}))={\bm{x}}(s_{i})$, and the related adjustments of the perturbation $\bm{\xi}(t_{i})\to\bm{\kappa}(s_{i})$.
All that is preventing one from readily following the steps in the pcLNA algorithm is the need for an explicit definition of the map G{G}. A natural question arises: does a universal definition of the map G{G} exist that would apply to all reaction networks? As we will soon demonstrate, the answer is no; the concept of phase is inherently linked to the dynamics of the system’s paths. Despite this, as will be discussed in the next section, the task of identifying G{G} can be significantly simplified by utilizing results from the theory of dynamical systems.
2.5 The phase correction map
We use three connected concepts from the theory of dynamical systems to guide our approach in defining the phase correction map. The first concept is topological equivalence, which allows us to group various systems exhibiting the same type of dynamics. The second concept is decomposition, enabling us to break down complex (potentially high-dimensional) systems into their (possibly low-dimensional) non-linear component and linear components. This reduction simplifies the challenges associated with studying these systems in general, and specifically the definition of the phase correction map. The third concept involves the existence of exemplar systems within each group that are well understood and can serve as useful references. In the following sections, we will elaborate on these three concepts and explain how we practically apply them in the construction of phase correction maps. We will then apply them to define a map that works for all systems presenting a Hopf bifurcation and a different map that works for all systems exhibiting bistability through a fold bifurcation.
2.5.1 Topological equivalence
Topological equivalence between two continuous dynamical systems of the form (6) implies that the dynamics of the two systems are qualitatively identical. This is ensured by showing the existence of a homeomorphism333a continuous bijective function between topological spaces linking their parameter spaces and an associated parameter-dependent homeomorphisms mapping orbits of one system to the orbits of the other system, preserving the direction of time. When topological equivalence occurs locally, i.e. in respective neighborhoods of the domains of the two systems, then we say that the two systems are locally topologically equivalent (l.t.e.) in those neighborhoods (see Definitions 2.14-5 in [34], or SI section S3).
A practical approach to identify l.t.e. between dynamical systems is to examine the eigenvalues of their Jacobian matrix near an equilibrium point. To expand on this, we introduce some terminology. Suppose that a system of the form (6) has an equilibrium point, $({\bm{x}}_{\alpha_{p}},\alpha_{p})$, i.e. $F({\bm{x}}_{\alpha_{p}},\alpha_{p})={\bm{0}}$. Let $_{0}, n+n_{+}$, and $n_{-}$ be the number of eigenvalues of the Jacobian matrix $\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})$ that have zero, positive, and negative real part. An equilibrium is hyperbolic if $n_{0}=0$ and non-hyperbolic otherwise. For two systems that each have a hyperbolic equilibrium, l.t.e. near their respective equilibria is secured if they have identical values of $n_{+}$ and $n_{-}$ (see Theorem 2.2 in [34]). Establishing l.t.e. near a non-hyperbolic equilibrium is more complicated but it can be achieved in some cases, as discussed below.
2.5.2 Decomposition
Near a non-hyperbolic equilibrium, all orbits of a system, as in (6), either lie on or converge (either as $t\to\infty$ or $t\to$-$\infty$) to an $n_{0}$-dimensional manifold, known as the centre manifold. This centre manifold is tangent to the hyperplane that contains all the eigenvectors corresponding to the $n_{0}$ eigenvalues with zero real part (see Theorem 5.1 and Lemma 5.1 in [34] or SI section S3).
Moreover, a system as in (6) with a non-hyperbolic equilibrium $({\bm{x}}_{\alpha_{p}},\alpha_{p})$ is l.t.e. to the following system;
$\left\{\begin{aligned} \dot{{\bm{u}}}&={\bm{A}}{\bm{u}}+{\bm{g}}({\bm{u}},{\bm{V}}_{\alpha}({\bm{u}})),\\ \dot{{\bm{v}}}&=-{\bm{v}},\\ \dot{\bm{w}}&=\bm{w},\end{aligned}\right$. (15)
near $({\bm{x}}_{\alpha_{p}},\alpha_{p})$. Here ${\bm{u}}\in\mathbb{R}^{n_{0}}$, ${\bm{v}}\in\mathbb{R}^{n_{+}}$, $\bm{w}\in\mathbb{R}^{n_{+}}$ are coordinates corresponding to the eigenbasis and $\bm{A}$} is an $n_{0}\times n_{0}$ matrix with all its eigenvalues, for $\alpha=\alpha_{p}$, having zero real part and ${\bm{g}}:\mathbb{R}^{n}\to\mathbb{R}^{n_{0}}$ is a smooth function with Taylor expansions starting with at least quadratic terms. The centre manifold is defined by ${\bm{V}}_{\alpha}({\bm{u}})$ in the sense that points $({\bm{u}},{\bm{v}},\bm{w})$ lie on the manifold if and only if they satisfy $({\bm{v}},\bm{w})={\bm{V}}_{\alpha}({\bm{u}})$ (see Theorem 5.4 in [34], [52] or SI section S3).
We can always compute the decomposition $\bm{J}({\bm{x}}_{\alpha},\alpha)={\bm{R}}^{T}{\bm{\Lambda}}{\bm{R}}$ where $\bm{\Lambda}$} the Jordan canonical form of $\bm{J}({\bm{x}}_{\alpha},\alpha)$ and specify coordinates $({\bm{u}},{\bm{v}},\bm{w})^{T}={\bm{R}}^{T}({\bm{x}}-{\bm{x}}_{\alpha})$. Then we can use the $n\times n_{0}$ matrix, say ${\bm{R}}_{1}$, with $n_{0}$ columns corresponding to the $n_{0}$ eigenvalues with zero real part when $\alpha=\alpha_{p}$ to derive a vector of $\bm{u}$}-coordinates of the orbits lying on the centre manifold. That is, take a long term orbit $\gamma=\{{\bm{x}}(t):t\geq 0\}$ near the equilibrium ${\bm{x}}_{\alpha}$ of the system in (6). The orbit $\gamma$ will either lie on or quickly converge to the centre manifold of (6) emerging near $({\bm{x}}_{\alpha_{p}},\alpha_{p})$. Then for ${\bm{x}}\in\gamma$, ${\bm{u}}={\bm{R}}_{1}^{T}({\bm{x}}-{\bm{x}}_{\alpha})$ is a $\bm{u}$} coordinate, as in (15), of a centre manifold satisfying an equation of the form $({\bm{v}},\bm{w})={\bm{V}}_{\alpha}({\bm{u}})$. This provides a practical way for finding $\bm{u}$}-coordinates for orbits lying on the centre manifold near a non-hyperbolic equilibrium of a system as in (6).
2.5.3 Normal forms
It is helpful, but not necessary, for the specification of the phase correction map of a class of l.t.e. systems, to identify a simple exemplar system in this class. We call normal forms the simplest systems in a class of l.t.e. systems. The normal forms of low dimensional centre manifolds are well studied. Characteristic examples include the fold, and Hopf bifurcations (see [34], and Figure 4).
Normal forms are useful in two ways. First, they specify the conditions under which a system is locally topologically equivalent (l.t.e.) to them. These conditions are often related to the eigen-decomposition of the Jacobian matrix in (7) at a non-hyperbolic equilibrium. This provides a practical, and often straightforward, method for determining whether a system belongs to a particular class. Secondly, their simpler form makes it easier to identify the key qualitative features (e.g., oscillations, multi-stability) shared by systems in that class, and to guide the choice of an appropriate phase correction map.
2.5.4 Consequences for phase correction
We proceed with the following actions based on the results obtained:
  1. 1.
    The concept of topological equivalence enables us to categorize different systems that exhibit qualitatively identical dynamics in the vicinity of an equilibrium point. We will employ the same form of phase correction map for each group of such systems.
  2. 2.
    The principle of decomposition allows us to construct a set of coordinates for a system where its non-linear dynamics are expressed. Given that phase drifts are relevant only to non-linear dynamics, we will apply phase correction to these particular coordinates.
  3. 3.
    The idea of normal forms helps us to determine whether a specific system belongs to a recognised class of systems with qualitatively identical dynamics near an equilibrium. Furthermore, it allows us to study the characteristics of these classes (e.g., oscillations, multi-stabilities) to select appropriate phase correction maps.
2.5.5 Generalised pcLNA
First, we make the observation that, while the Algorithm 1 uses a single deterministic trajectory $\gamma=\{{\bm{x}}(t):t\geq 0\}$, some classes of systems prompt the need for using more than one trajectory. This is particularly the case for systems presenting more than one stable equilibria in nearby locations (see Figure 2D and section 4). The generalised pcLNA algorithm in 2 allows for a phase correction function that maps a stochastic state $\bm{X}$} to a point lying on one of a set of deterministic trajectories $\gamma^{(j)}=\{{\bm{x}}^{(j)}(t):t\geq 0\}$, $j=1,2,\dots,J$, that all satisfy the system in (6) and differ only on their initial conditions, ${\bm{x}}^{(j)}(0)$. It is clear that Algorithm 1 is a special case of Algorithm 2 for $J=1$.
Algorithm 2 Generalised pcLNA algorithm
Inputs:
a. Initial conditions: $s_{0}=0=t_{0}$, $j_{0}$, ${\bm{X}}(0)$, ${\bm{x}}^{(j)}(0)$, $j=1,\dots,J$. Parameters: $\Delta t$, $\Omega$.
b. Solutions $\gamma^{(j)}=\{{\bm{x}}^{(j)}(t):t\geq 0\}$, $j=1,2,\dots,J$ of a RRE as in (6) and the corresponding solutions ${\bm{C}}^{(j)}$ and ${\bm{D}}^{(j)}$ of (12), (13), respectively.
c. The phase correction map $G:\mathbb{R}^{n}\to\{\gamma^{(j)}:j=1,\dots,J\}$
Steps: Take $\bm{\kappa}(s_{0})=\Omega^{1/2}({\bm{X}}(0)-{\bm{x}}^{(j_{0})}(0))$ and for $i=1,2,\dots$,
  1. 1.

### v4

@@ 2.2 System Size Expansion
Following the work of [35] and [66], it is common in seeking methods to approximate the process $\{{\bm{Y}}(t):t\geq 0\}$, to introduce a system size parameter $\Omega$. Through setting ${\bm{X}}(t):={\bm{Y}}(t)/\Omega$, one can study the dependence of stochastic fluctuations upon system size. It is sufficient to assume that the rates $\pi_{j}({\bm{Y}}(t))$ depend upon $\Omega$ as $\pi_{j}({\bm{Y}}(t)):=\Omega\rho_{j}({\bm{Y}}(t)/\Omega)$ where $\rho_{j}:\mathbb{R}^{n}\to[0,\infty)$ are called macroscopic rates of the system and are smooth. This is a fairly-relaxed condition, that can be further relaxed to more general relations between the two functions [66, 34].
@@
${\bm{X}}(t+dt)-{\bm{X}}(t)=\sum_{j=1}^{r}\frac{{\bm{\nu}}_{j}}{\Omega}N_{j}\left(\Omega\rho_{j}({\bm{X}}(t))dt\right),\quad dt>0$ (5)
with ${\bm{X}}(0)={\bm{X}}_{0}$, thus providing a time evolution law for the process $\{{\bm{X}}(t):t\geq 0\}$. If, for each $t\geq 0$, we define ${\bm{x}}(t)$ as the limit in probability of ${\bm{X}}(t)$ as $\Omega\to\infty$, we can use the law of large numbers  [^3] Note that for a sequence $X_{1},X_{2},\dots$, of independent Poisson random variables with mean $0<\lambda<\infty$, $N^{-1}\sum_{i=1}^{N}X_{i}$ converges in probability to $\lambda$, as $N\to\infty$. (see [35, 34]) to derive the limit of equation (5), as $\Omega\to\infty$,
@@
Letting $dt\to 0$ we derive the macroscopic reaction rate equation (RRE), corresponding to the classical mass-action kinetics. Typical in applied biochemical settings is the need to study the qualitative changes in the dynamics of $\{{\bm{X}}(t):t\geq 0\}$ and $\{{\bm{x}}(t):t\geq 0\}$ when one (or more) of the system parameters – typically reaction rate constants – are allowed to vary smoothly. Let this parameter be given by $\alpha\in\mathbb{R}$ and write the RRE as
$\dot{{\bm{x}}}=\frac{d{\bm{x}}}{dt}={\bm{F}}({\bm{x}},\alpha),\quad{\bm{F}}({\bm{x}},\alpha):=\sum_{j=1}^{r}{\bm{\nu}}_{j}\rho_{j}({\bm{x}}(t),\alpha),\qquad{\bm{x}}(0)={\bm{x}}_{0}.$ (6)
where both $\rho_{j}$ and $F$ depend smoothly on $\bm{x}$} and $\alpha$. The corresponding Jacobian matrix is
@@
Example (SIR model, continued)
@@ 2.3 The Linear Noise Approximation (LNA)
The time-evolution of $\{\bm{\xi}(t),t\geq 0\}$, for sufficiently large $\Omega$, can be described (see [49] or SI, section S2 for the derivation details) by the linear Stochastic Differential equation (SDE) in the Itô sense
@@
where $\bm{J}_{t}=\bm{J}({\bm{x}}(t))$ is the Jacobian matrix in (7) evaluated at ${\bm{x}}(t)$, ${\bm{B}}_{t}$ an $r$-dimensional Wiener process [72], and
@@
the product of the stoichiometry matrix ${\bm{S}}=[{\bm{\nu}}_{1}\cdots{\bm{\nu}}_{r}]$, and the square root of the diagonal matrix, with main diagonal entries the reaction rates $(\rho_{1}({\bm{x}}(t)),\dots,\rho_{r}({\bm{x}}(t)))$. This linear SDE in (9) has a solution that, for $0\leq s\leq t<\infty$, can be written as
$\bm{\xi}(t)={\bm{C}}(s,t)\bm{\xi}(s)+{\bm{\eta}}(s,t),\quad{\bm{\eta}}(s,t)\sim\text{MVN}({\bm{0}},{\bm{D}}(s,t)),$ (11)
@@
The above representation implies that by solving the initial value problems in (6), (12) and (13), and starting with an initial condition for $\bm{\xi}(0)=\bm{\xi}_{0}$ drawn from an arbitrary MVN distribution, one can easily derive the MVN probability distribution of $\bm{\xi}(t)$, for any time $t>0$. In this case, we say that the LNA prediction for ${\bm{X}}(t)$ was initialised at time $s=0$. The process $\{{\bm{X}}(t):t\geq 0\}$ modeled by the LNA is only an approximation of the true dynamics of the reaction network $\{{\bm{Y}}(t)/\Omega:t\geq 0\}$ where $\{{\bm{Y}}(t):t\geq 0\}$ evolves according to (4).
Therefore, when we refer to the notion of accuracy, we mean the accuracy in the distribution of ${\bm{X}}(t)$ when trying to capture the distribution of ${\bm{Y}}(t)/\Omega$ for each $t\geq 0$, where $\{{\bm{Y}}(t):t\geq 0\}$ evolves according to (4).
@@
These matrices depend on the solution, $\{{\bm{x}}(t)=(x_{1}(t),x_{2}(t),x_{3}(t))^{\top}:t\geq 0\}$, of the RRE in (6). Solving (12) and (13) using the above matrices gives the matrices ${\bm{C}}(s,t)$ and ${\bm{V}}(s,t)$ required to derive the distribution of $\bm{\xi}(t)$ in (11) and in turn the state ${\bm{X}}(t)$ in (8).
@@ 2.3.1 Accuracy of the LNA
The LNA accurately approximates the master equation in specific contexts. As discussed in SI (see section S2), the LNA is derived using the Central Limit Theorem, as $\Omega$ tends to $\infty$, as an approximation of the probability distribution function satisfying the master equation. As a result, we expect the LNA to be accurate when $\Omega$ is sufficiently large. The value of $\Omega$ needed to ensure accuracy depends on the characteristics of the system’s dynamics and the length of time internal where the approximation is sought.
For any given finite $\Omega$, if the LNA accurately represents the distribution of the state of a given system at some time-point, it is expected to maintain this accuracy for a subsequent short time interval (see [69]). The duration for which the LNA remains accurate is influenced by the dynamics of the system and the level of stochasticity present. For instance, it was observed in [46] that for periodic systems with a moderate size $\Omega$, the time interval over which the LNA retains its accuracy is approximately equal to the period of the system.
@@
It is important to note that, for moderate system sizes, the LNA tends to fail in approximating the long-term dynamics of multi-stable and oscillatory systems (see [57, 5, 29, 64, 46] and Figure 2(C-D)). We further discuss the failure of the LNA in Section 3.1.
3 The phase corrected Linear Noise Approximation (pcLNA)
In this section, we develop the phase corrected Linear Noise Approximation (pcLNA) algorithm for stochastic simulation of population dynamics. We begin by introducing a general definition of phase.
3.1 Phase
Important in diagnosing this failure of the LNA approximation, and indeed the rest of this paper, is the idea of phase. Existing definitions of phase for stochastic systems typically extend the concept of phase for noisy limit cycles [56, 6, 11]. Here we introduce a definition suitable for any non-linear dynamics arising from the RRE (6), including systems with oscillations or multistability.
We use the term phase in its common-language sense: a stage in a dynamic process. To define the phase of an arbitrary point $\bm{X}$} on a stochastic trajectory $\{{\bm{X}}(t):t\geq 0\}$, we compare it with a set of reference trajectories
$\Gamma=\bigcup_{\ell=1}^{L}\{{\bm{x}}^{(\ell)}(t):t\geq 0\},\qquad L\geq 1,$ (14)
generated from the system in (6). The choice of $\Gamma$ depends on the underlying dynamics and will be discussed in Section 3.3. For example, if the deterministic system has an attracting limit cycle, an appropriate choice is a single trajectory lying on that cycle; if it has multiple stable equilibria, $\Gamma$ may consist of trajectories converging to each equilibrium.
We find an appropriate metric $d$ to measure the distance between $\bm{X}$} and points on the trajectories in $\Gamma$. Based on metric $d$, we identify the point $G({\bm{X}})\in\Gamma$ that is closest to $\bm{X}$}. The phase of $\bm{X}$} is the time coordinate $s\geq 0$ of this closest point. Formally, the phase $s$ of $\bm{X}$} is defined via
$(j,s)=\operatorname*{argmin}_{\ell\in\mathcal{L},\,t\geq 0}d({\bm{X}},{\bm{x}}^{(\ell)}(t)),\qquad\mathcal{L}=\{1,\dots,L\},$ (15)
and the associated reference point is obtained through the map $G:\mathbb{R}^{n}\to\Gamma$,
$G({\bm{X}})={\bm{x}}^{(j)}(s).$ (16)
The most appropriate metric $d$ to be used for the phase definition above depends on the system dynamics and will be discussed in Section 3.3. Aside from the phase-defining metric $d$ and the reference trajectories $\Gamma$ that are specific to the system dynamics, this definition of phase applies to any type of dynamics, not only oscillatory or multistable systems.
While the LNA is accurate for systems with linear dynamics (see Section 2.3.1), for systems with non-linear dynamics (e.g. oscillatory or multi-stable), as $t$ increases the phase of ${\bm{X}}(t)$ will likely be very different to $t$. This means, with high probability, there exists some time point $\tau>0$ such that ${\bm{X}}(\tau)$ is at a very different stage of its dynamic process to the deterministic ${\bm{x}}(\tau)$ despite ${\bm{X}}(0)={\bm{x}}(0)$. Since the drift and diffusion in equation (9) which control the dynamics of ${\bm{X}}(t)$ for $t\geq\tau$ under the LNA depends only on ${\bm{x}}(t)$, the LNA will make predictions of ${\bm{X}}(t)$ for $t\geq\tau$ as if its in the same stage of its dynamic process as ${\bm{x}}(t)$ rather than considering the location of ${\bm{X}}(t)$. This leads to the LNA failing to capture the SSA dynamics of the reaction network $\{{\bm{Y}}(t)/\Omega:t\geq 0\}$ where $\{{\bm{Y}}(t):t\geq 0\}$ evolves according to (4). We call this phenomenon a phase drift.
@@ 3.1 Phase
Therefore, initialising the LNA to the correct phase is a necessary condition for the LNA to be accurate for non-linear systems. This leads us to speculate that, if one can continually initialise the LNA with the correct phase so that the phase of the stochastic and deterministic trajectories remain in agreement, then the LNA will remain accurate at all times and produce paths agreeing with those from the SSA. We argue that, for some classes of dynamical systems, this is indeed the case. For instance, it is shown in [46] that this is true for systems whose RRE has an attractive limit cycle solution and a range of moderate system sizes.
This is the basis of the simulation algorithm developed next. We adopt the term used in [46] to name this simulation algorithm, phase corrected Linear Noise Approximation (pcLNA).
Figure 2: LNA method summary. A. LNA is a stochastic approximation of the stochastic population dynamics that can be exactly simulated using the SSA. LNA uses the numerical solutions of the deterministic RRE and a SDE that can be solved exactly. The SIR population dynamics model is used here as an example. B. Comparison between SSA and LNA on the SIR population dynamics in terms of simulated $99\%$ Confidence Intervals (CI) shows astonishing agreement as the blue lines and shades almost completely overlap with the pink lines and shades
. C. In contrast, same comparisons between the LNA and SSA in one of the three species of the oscillatory system in [71] show significant discrepancies. D. Significant discrepancies are also observed for the same comparisons on the bistable Genetic Toggle-Switch system in [14].
3.2 The pcLNA algorithm
@@
Here ${G}:\mathbb{R}^{n}\to\Gamma$, is the phase-correcting function (16) that maps any stochastic state ${\bm{X}}(t)\in\mathbb{R}^{n}$ to a point lying on a solution in some set of reference RRE solutions $\Gamma=\bigcup_{\ell=1}^{L}\{{\bm{x}}^{(\ell)}(t):t\geq 0\}$ with its corresponding time being the phase $s\geq 0$ of the stochastic state ${\bm{X}}(t)$. The term $\bm{\kappa}(t)$ denotes the deviation of ${\bm{X}}(t)$ away from ${G}({\bm{X}}(t))$ scaled by $\Omega^{-1/2}$. We use this ansatz to simulate stochastic trajectories as described in Algorithm 1.
@@
a. Initial conditions: $t_{0}=0$, ${\bm{X}}(0)$, ${\bm{x}}^{(\ell)}(0)$, $\ell=1,\dots,L$. Parameters: $\Delta t$, $\Omega$.
b. Solutions $\Gamma=\bigcup_{\ell=1}^{L}\{{\bm{x}}^{(\ell)}(t):t\geq 0\}$ of a RRE as in (6) and the corresponding solutions ${\bm{C}}^{(\ell)}$ and ${\bm{D}}^{(\ell)}$ of (12), (13), respectively.
c. A metric $d$ on $\mathbb{R}^{n}$ for the phase-correcting map $G:\mathbb{R}^{n}\to\Gamma$ as in (15), (16).
Steps: Set $(j_{0},s_{0})$ such that ${G}({\bm{X}}(0))={\bm{x}}^{(j_{0})}(s_{0})$ and take $\bm{\kappa}(s_{0})=\Omega^{1/2}({\bm{X}}(0)-x^{(j_{0})}(s_{0}))$. Then, for $i=1,2,\dots$,
@@
    draw sample $\bm{\xi}(s_{i-1}+\Delta t)$ from the MVN distribution with mean ${\bm{C}}^{(j_{i-1})}(s_{i-1},s_{i-1}+\Delta t)\bm{\kappa}(s_{i-1})$, and variance matrix ${\bm{D}}^{(j_{i-1})}(s_{i-1},s_{i-1}+\Delta t)$;